OpenStudy (ashtorah13):

What is the next number in the sequence: 9, 3, 1, 1/3,... Answer is 1/9 but i am unsure how they got that answer, can someone show me how they got it? I get lost after getting 1/3, I thought it decreased by dividing all the numbers by 3 but that would mean that the answer should be .11111111 repeating not 1/9?

ganeshie8 (ganeshie8):

Hint : \[0.111\ldots = \dfrac{1}{10} + \dfrac{1}{100}+\dfrac{1}{1000}+\cdots \]

OpenStudy (ashtorah13):

Hahahaha I feel dumb 9/1 3/1 1/1 1/3 1/3 They are mirroring ?

OpenStudy (ashtorah13):

That was suposed to be 1/9 the last one...

ganeshie8 (ganeshie8):

your teacher must be feeling patterny :)

ganeshie8 (ganeshie8):

did you get my earlier hint ?

OpenStudy (ashtorah13):

Im studying for entrance exams. I want to go to collage.

OpenStudy (ashtorah13):

I thought I did but no, I didnt.

ganeshie8 (ganeshie8):

Ohkay, lets take it step by step and slow

ganeshie8 (ganeshie8):

look at the number \[246.0\] what does it represent ?

ganeshie8 (ganeshie8):

it is a shortcut for below lengthy expression : \[2\times 10^2 + 4\times 10^1 + 6\times 10^0 + 0\times 10^{-1}\] yes ?

OpenStudy (ashtorah13):

Sorry I am usually not this slow. I would say 246/1 right? and I don't understand your last comment.

ganeshie8 (ganeshie8):

how do you "spell out" the number 246.0 ?

OpenStudy (ashtorah13):

Two hundred fourty six

ganeshie8 (ganeshie8):

good, let me break that : ``` two hundred fourty six ```

ganeshie8 (ganeshie8):

same as ``` 200 + 40 + 6 ```

OpenStudy (ashtorah13):

oh ok

ganeshie8 (ganeshie8):

same as ``` 2x100 + 4x10 + 6x1 ```

OpenStudy (ashtorah13):

Now I understand your one comment.

ganeshie8 (ganeshie8):

246 represents : 2 hundreds 4 tens 6 ones

ganeshie8 (ganeshie8):

lets do a quick example

ganeshie8 (ganeshie8):

"spell out" the number 4789

OpenStudy (ashtorah13):

Four thousand seven hundred eighty nine

ganeshie8 (ganeshie8):

can you write it in numbers...

OpenStudy (ashtorah13):

4000 700 80 6

ganeshie8 (ganeshie8):

``` 4 x 1000 + 7 x 100 + 8 x 10 + 9 x 1 ```

OpenStudy (ashtorah13):

9

ganeshie8 (ganeshie8):

can we write it as : ``` 4 x 10^3 + 7 x 10^2 + 8 x 10^1 + 9 x 1^0 ``` ?

OpenStudy (ashtorah13):

I put a 6 not a 9. Continue sorry.

ganeshie8 (ganeshie8):

above is called "decimal expansion" of 4789

ganeshie8 (ganeshie8):

See if you can write the decimal expansion of 345

OpenStudy (ashtorah13):

3x10^2 + 4x10^1 + 5x1^0

ganeshie8 (ganeshie8):

Excellent! lets do one more example

ganeshie8 (ganeshie8):

write the decimal expansion of 328.6

OpenStudy (ashtorah13):

3x10^2 + 2x10^1 + 8x10^0 + .6X.1^0

ganeshie8 (ganeshie8):

wrong 2, 1, 0, what comes next ?

OpenStudy (ashtorah13):

.1

OpenStudy (ashtorah13):

-1

ganeshie8 (ganeshie8):

-1, yes

ganeshie8 (ganeshie8):

so the decimal expansion of 328.6 is ``` 3x10^2 + 2x10^1 + 8x10^0 + 6x10^(-1) ```

OpenStudy (ashtorah13):

why would it be a negative instead of a decimal?

ganeshie8 (ganeshie8):

that is a very good question! before answering that, could you "spell out" the number 328.6 ?

OpenStudy (ashtorah13):

three hundred twenty eight and six tenths

ganeshie8 (ganeshie8):

very good how do you "write out" six tenths ?

ganeshie8 (ganeshie8):

i mean, how do you express six tenths using numbers ?

OpenStudy (ashtorah13):

6x10^(-1)

ganeshie8 (ganeshie8):

how did u get -1 ?

ganeshie8 (ganeshie8):

you're correct btw..

OpenStudy (ashtorah13):

Thats what you showed me when we expressed 328.6

OpenStudy (ashtorah13):

I don't know how, that's what I was shown.

OpenStudy (freckles):

do you know six-tenths can also be written as 6/10? @Ashtorah13

OpenStudy (ashtorah13):

Yes

OpenStudy (freckles):

\[\frac{6}{10}=\frac{6}{10^{1}}=6 \cdot \frac{1}{10^{1}}=6 \cdot 10^{-1}\]

OpenStudy (ashtorah13):

Still don't know why its a negative and not a decimal.

OpenStudy (freckles):

the exponent is negative not the number itself the exponent is negative because it is a decimal

OpenStudy (ashtorah13):

Ok so its just a different way to show the decimal?

OpenStudy (freckles):

examples: \[.6=\frac{6}{10} \text{ or } 6 \cdot 10^{-1} \\ .06 =\frac{6}{100}=\frac{6}{10^2} \text{ or } 6 \cdot 10^{-2} \\ .006 =\frac{6}{1000}=\frac{6}{10^3} \text{ or } 6 \cdot 10^{-3}\]

OpenStudy (ashtorah13):

OH ok so it shows where the decimal will go!

OpenStudy (ashtorah13):

so like 6X10^(-5) would be .00006

OpenStudy (freckles):

yes like 6*10^2 would be 06.00 move the decimal over 2 spaces to the right so you have 0600. or just 600 but 6*10^(-2) would be 06.00 move the decimal over 2 spaces to left so you have .0600 or just .06

OpenStudy (freckles):

yes

OpenStudy (ashtorah13):

Bingo Thank you. But I still don't understand my original question...

OpenStudy (freckles):

what is the original question

OpenStudy (freckles):

are you talking about the next number in that sequence you first mentioned?

OpenStudy (ashtorah13):

Yes

OpenStudy (ashtorah13):

I feel what I just learned ties into it but I am unsure

OpenStudy (freckles):

do you know what common ratio means?

OpenStudy (ashtorah13):

No? (I may, I just don't remember.)

OpenStudy (freckles):

oh actually you did see the common ratio earlier what you meant to say earlier is that the patterned looked like all you had to do was take term and divided by 3 to get next term right?

OpenStudy (ashtorah13):

Ya

OpenStudy (freckles):

another way to say divide by 3 is to say multiply by 1/3

OpenStudy (freckles):

9*1/3=3 3*1/3=1 1*1/3=1/3 1/3*1/3=?

OpenStudy (ashtorah13):

It makes perfect since now!

OpenStudy (freckles):

don't whip out your calculator

OpenStudy (freckles):

just multiply straight across :)

OpenStudy (freckles):

and you are right .11111111111111111111111111111111111111111111111111111..... is 1/9 :)

OpenStudy (freckles):

most calculators won't tell you that though a bad side effect of some calculators :p

OpenStudy (ashtorah13):

common numerator dont need to, :3 I knew it was 1/9 just wasnt sure how they got it. and oh my, how the heck would you simplify .1111111111 into 1/9?

OpenStudy (ashtorah13):

Is there a way to give you both medals?

OpenStudy (freckles):

we can even prove 1/9 is .1111111111111111111111111111111111111111(repeating) \[x=.1111111111111111111.... \\ \text{ multiply both sides by 10 } \\ 10x=1.111111111111111111111................... \\ \text{ now take equation 2 and subtract from \it equation 1 } \\ 10x-x=1.111111111111111111111...........-.1111111111111111111.... \\ 9x=1.00000000000000000..... \\ 9x=1 \\ \text{ divide both sides by 9 } \\ x=\frac{1}{9}\]

OpenStudy (freckles):

therefore .111111(repeating)=1/9

OpenStudy (freckles):

another example: Say we wanted to figure out what fraction .534534534534.... equals let x=.534534534534.... multiply both sides by 1000 repetition begins after 1000th place \[1000x=534.534534534......\] now we subtract 2nd from 1st equation again \[999x=534.00000000000000000000.... \\ \text{ so } x=\frac{534}{999}\] check your calculator if you want to

OpenStudy (freckles):

you will see 534/999 is .534(repeating)

OpenStudy (freckles):

now we subtract 1st from 2nd equation *

OpenStudy (ashtorah13):

Is anything repeating over 9? Say you have .7777777 would that mean it would be 777/999?

OpenStudy (freckles):

let's find out let x=.7777777777777(repeating) repetition starts after 10th's place so we multiply both sides by 10 10x=7.77777777777(repeating) subtract 1st from 2nd we get 10x-x=7 9x=7 solve for x by dividing both sides by 9 x=7/9 and you said 777/999 which still works since 777/999 can be reduced to 7/9 but yeah you are right (you just have to put the right amount of 9's depending on the repetition) sometimes you can reduce the answer just like what you found with 777/999

OpenStudy (freckles):

another example for fun \[x=.3401340134013401.... \\ \text{ here we multiply both sides by } 10000 \\ \text{ since repetition begins after } 10000th \text{ place} \\ \\ 10000x=3401.34013401...... \\ \text{ subtract 1st from 2nd } \\ 10000x-x=3401 \\ 9999x=3401 \\ x=\frac{3401}{9999}\]

OpenStudy (ashtorah13):

10x=7.77777777777 - x=.7777777777777 (let me see if I can explainthis right) I was uncertine how you kept getting 9 from 10 after subtracting. Is it when you subtract 10x from x you have to put a 1 infront of the x? Making it 10x- 1x? So you get 9x?

OpenStudy (freckles):

that is correct x is 1x

OpenStudy (ashtorah13):

Oh my this is making more and more since. My algebra teacher wasn't vary good.

OpenStudy (freckles):